MARS DESERT RESEARCH STATION

EVA Report – April 15th

Crew 178 EVA #5 report 15 APR 2017

Weather: sunny, dry and hot
Temperature: around 30°C
Location: 12 S 0519938 4247631
Elevation: 1333 m
Duration: 120 min
Team: Damien Mertens (crew engineer), Elke Mergny (crew geologist), Quentin Thomas (crew astronomer)
Purpose:
–       Geophysical survey of the area with a ground penetrating radar
–       Telecommunication: installation of marking devices that use tools to cartography the sound beacons placed next to dangerous reliefs and hence, to construct a map of risk areas.
Outputs:
–       Ground penetrating radar: 7 geophysical profiles collected (area of the survey=10000m²). 3D mapping will be established
–       The installation of the marking device succeeded but the localisation system needs still to be adapted (see tomorrow’s EVA)

Crew Photos – April 15th

Careful driver

 

Checking the zone

 

Spacesuit room

 

Stay serious

 

This is the hab

 

Bread or pizza

Journalist Report – April 15th

Crew 178 Journalist report Sol 5  15April2017
Prepared by: Mathieu « Mitch » Vander Donckt, Crew 178 Journalist and Scientist
Crew 178 – UCL to Mars
It is astonishing how the situation can deteriorate in a few hours. Being isolated, we cannot depend on external assistance. We need to rely on our own skills and training and have a blind trust in the abilities of our crewmates. That was proved during last day’s incidents, that we afterward named “the cowboy crisis”.
Sol 3 ended with the disappearance of two crew members. The dust storm that raged that night made any expedition to lead an emergency rescue, impossible. Besides, protocol forbids to go outside of the base at night, whatever happens. It is of course for our own safety, the lives of all the crew cannot be put at risk for an operation with such uncertain outcomes. However, it was thwarting to stay inside in such a time. We defined an area of search near the sector they were surveying the last time we had contact with them and decided to send one of Tarzan’s probes with an infrared camera. Due to the storm, it crashed into the Northern Rim, a mountain range north of the station. It was a risk to take, and worth it! On the last images send by the probe, we could see our two fellows sheltered at the base of the Rim.
A few hours and some minor incidents later, they were in the hands of “Coach”, which is not only our persecutor – I often hear cries of pain from the first floor when she “wants to make us more fit and healthy than when we were on Earth” with her workout – but is also our Health and Safety Officer. Tonight, we will use our only rations of fresh food to celebrate their return and the halfway through our stay on Mars. Tarzan and Patch told us how, after a weary night without sleep, they encountered cowboys. As you know, that’s impossible on Mars. I guess it was tiredness, the low visibility, and queer-shaped rocks. Or maybe they have gone crazy… We will see that in the following days.

Sol Summary – April 15th

Crew 178 Daily Summary Report 15APR2017

 

MDRS Daily Summary Report for sol 5

 

Summary Title: Fifth EVA

 

Mission Status:  Good

 

Sol Activity Summary:

 

Hab

8.30:   Breakfast

11.00: Muscle wasting experiment: exercise session

12.30: Lunch

 

GreenHab

/

 

ScienceDome

14.00: End setup recycling experiment. 

 

EVA

14.00 UTM NAD27 12S 4247631-519930

Ground penetrating radar experiment : data collection

Topography Geomatic study: fixing probe

Mapping of dangerous areas with the use of probes

 

Weather:

Sunny

 

Crew Physical Status:

Everyone is fine

 

Look Ahead Plan:

EVA + Experiments

 

 

Reports to be file: Operation Report , Daily Summary Report, HSO Report, Journalist Report

 

Support Requested:

Information about water consumption

Bottom part of the hab outside repair

What sould we do with the inside toilet ?